Marcus is 6 yaers older then Max, the sum of their age is 38, how old is Marcus and how old is Max
Answer by ikleyn(52946)

.
Marcus    + Max = 38


(Max + 6) + Max = 38


2*Max = 38 - 6 = 32


  Max = 32/2 = 16 years old.


Marcus = Max + 6 = 16 + 6 = 22 years old.

Solved.